# SMooGPT: Stylized Motion Generation using Large Language Models

### v1

@@ 5. Results and Evaluation
SMooDi (in-domain) (Zhong et al., 2025) 71.698 4.727 0.528 0.172
SMooDiw/o-CG(in-domain) (Zhong et al., 2025) 25.471 4.316 0.597 0.135
SMooGPT (Ours) (in-domain) 44.968 4.390 0.610 0.106
SMooDi (out-domain) (Zhong et al., 2025) 36.792 4.172 0.616 0.182
SMooDiw/o-CG(out-domain) (Zhong et al., 2025) 23.584 3.898 0.626 0.132
SMooGPT (Ours) (out-domain) 40.174 4.163 0.551 0.124
Evaluation metrics.
We evaluate our method along three dimensions: content preservation, style reflection, and realism. For the assessment of content preservation, we employ metrics used in (Chen et al., 2023): motion-retrieval precision (R-Precision), Multi-modal Distance (MM Dist). The two metrics are calculated on the HumanML3D test set. Additionally, we exploit the foot skating ratio (FS-Ratio) proposed by (Karunratanakul et al., 2023) for our motion realism evaluation. For style reflection, we employ Style Recognition Accuracy (SRA) (Jang et al., 2022), calculated against the motion in the 100STYLE test set. We do not include the common FID metric for evaluating content preservation due to its inherent ambiguity in our task. See supplementary for the detailed discussion.
5.1. Comparison
To comprehensively demonstrate the performance of our method, we conduct comparisons across two subtasks in what follows.
Text-guided Stylization.
This task generates stylized motion based on both the content and style texts. Two concurrent work StyleMotif (Guo et al., 2025) and MulSMo (Li et al., 2024b) can produce stylized motion in this setting, however, their code is not publicly available at the time of submission. We thus create two baseline methods using a pre-trained LLM (i.e., ChatGPT-3.5 (OpenAI, 2024)) and two motion generators (i.e., MotionGPT (Jiang et al., 2023) and MLD (Chen et al., 2023)). Strategically, we first ask ChatGPT to combine the content and style texts, and employ the motion generator to produce the corresponding motion according to the composed text. Statistical results are displayed in Tab. 1(top). In terms of style reflection (SRA), our approach outperforms two baseline methods by an order of magnitude (i.e., increased by $338\%$ and $591\%$, respectively), at an average cost of 10.1% sacrifice (i.e., 14.2% on MM-Dist, 5.5% on R-Precision, and 10.7% on FS-Ratio) on motion realism and content preservation, setting a new state-of-the-art for the text-guided stylization. The relatively small drop indicates the trade-off between style reflection, content preservation, and motion realism  (Zhong et al., 2024): stronger style expression may naturally cause partial drift from the original content text, slightly lowering content preservation and motion realism. The corresponding visual results are presented in Fig. 6(top). In two typical cases, baseline methods either cannot preserve the content or fail to reflect the style.
Motion-guided Stylization
This task generates stylized motion based on the content text and the style motion sequence. To the best of our knowledge, SMooDi (Zhong et al., 2024) is the only work for this purpose. Consequently, we focus on comparing our approach with SMooDi in this setting. To ensure a comprehensive and fair comparison with SMooDi, we fine-tune our model on the 100STYLE dataset and apply a few additional modifications. Specifically, SMooDi relies on classifier-based style guidance derived from a style encoder trained on the 100STYLE dataset. However, because the encoder is limited to a small set of locomotion-focused styles, its guidance works well for seen styles but degrades for unseen ones. To evaluate generalization, we therefore divide both the training and evaluation sets into in-domain and out-of-domain subsets (see supplementary). In addition to the original SMooDi, we also compare against a variant without classifier guidance, denoted as SMooDiw/o-CG, to further analyze the impact of the guidance mechanism.
The statistical and visual comparisons are presented in Tab. 1 (bottom) and Fig. 6 (bottom), respectively. Quantitatively, under the in-domain setting, our method is worse than SMooDi in terms of style reflection. This is primarily because classifier-based guidance provides more accurate direction for seen styles. However, relying on classifier-based guidance has two notable limitations. First, it requires a style motion sequence at inference time, making it incompatible with text-driven stylization. Second, because the classifier-based guidance is provided by a style encoder trained on the 100STYLE dataset, it implicitly biases the model toward locomotion-related content. Therefore, compared with SMooDiw/o-CG, we outperform it on SRA by a noticeable 76.5%. In the out-of-domain setting, our method outperforms SMooDi both with and without classifier-based guidance, achieving a 9.2% and 70.3% improvement in SRA, respectively. When transitioning from the in-domain to the out-of-domain setting, SMooGPT shows a relatively small performance drop of 10.7% in style-related accuracy, while SMooDi suffers a substantial decline of 48.7%. This highlights the better generalization ability of our method to new styles.
@@ 5.1. Comparison
Method SRA $\uparrow$ MM Dist $\downarrow$ R-P.(Top-3) $\uparrow$ FS-Ratio $\downarrow$
w/o BP-Space 17.96 5.212 0.524 0.103
w/o Pre-train 25.15 5.190 0.541 0.136
Ours 33.33 5.096 0.563 0.141
@@ 5.2. Ablation Study
We conduct several ablation experiments on text-guided stylization to validate the effectiveness of our framework design choices. We summarize key findings below.
The importance of body-part Space
To assess the impact of incorporating the body-part text space, we conduct an ablation study using the original MotionGPT (Jiang et al., 2023) model—referred to as w/o BP-Space—which is trained solely on global textual inputs within the same dataset. As illustrated in Tab 2, our method achieves a two times improvement in SRA, underscoring the efficacy of integrating body-part information. While metrics related to motion realism and content preservation exhibit only minor declines when body-part texts are excluded, the substantial gain in SRA highlights the advantage of a structured representation, which facilitates fine-grained control over individual body parts, leading to enhanced alignment between textual descriptions and generated motions
The importance of Pre-training
To assess the role of pretraining, we conduct an ablation study by removing the pretraining phase (denoted as w/o Pretrain). As presented in Tab. 2, the absence of pretraining leads to a decline in all evaluation metrics, with the SRA experiencing a significant drop of up to 25%. This performance degradation underscores the importance of the pretraining phase, which serves to align body-part texts with their corresponding motion representations. Such alignment establishes a coherent latent space, facilitating more effective and efficient fine-tuning during the subsequent post-training stage, ultimately enhancing overall model performance.
@@ 5.3. User Evaluation
Figure 7. User Evaluation Statistics. The percentage of times our approach is preferred over (a) ChatGPT+MLD and (b) ChatGPT+MotionGPT is reported. When choosing from a pair of generated motions, users are asked to evaluate three aspects - Content Preservation, Style Reflection, and Motion Realism. The higher the percentage (50% is a tie), the better our results.
We conduct a user perceptual study using pairwise comparisons on the text-guided stylization task. In each test case, participants are presented with a motion content text, a motion style text and a pair of motion clips - one produced by our SMooGPT and the other by a competing method (i.e., ChatGPT+MLD and ChatGPT+MotionGPT). For each pair of motions, participants are asked to evaluate three aspects - content preservation, style reflection, and motion realism, as shown in the legend in Fig. 7. In total, we have invited $50$ participants, and the percentage of times our approach is preferred over the competing methods is displayed in Fig. 7. Our SMooGPT performs favorably against ChatGPT+MotionGPT over all three aspects, with $81.3\%$, $81.7\%$, and $80.7\%$, respectively. Similarly, compared with ChatGPT+MLD, although the percentage our method wins drops around $7\%$, it is still considerably high (over $72\%$). Overall, the user perceptual study validates our superior performance and is consistent with observations in the comparison.
@@ 5.4. Discussions
Fueled by the strong reasoning capabilities of LLMs, our approach serves as an effective composer that can reconcile motion content and style texts in a conflict-free manner. Resolving such conflicts is crucial for stylized motion generation, as content and style often impose contradictory constraints on specific body parts. Fig. 8 illustrates a representative example: the motion content ”throw” suggests that one arm should stay close to the body, whereas the motion style ”Aeroplane” requires both arms to be lifted and extended outward. Our method successfully resolves this conflict by selectively borrowing one arm configuration from each source—resulting in a composed motion where one arm performs the throwing action, while the other adopts the aeroplane pose. In contrast, SMooDi, which takes the style motion (rather than style text) as input, fails to produce a motion that coherently reflects both the content and the style.
Style and body-part alignment capability
To assess whether LLMs can understand the relationship between styles and the body parts they affect, we conducted a quantitative evaluation on the 100STYLE dataset. Specifically, we prompted GPT to predict the most relevant body parts for each style using only its global style name and description, and compared its predictions to human-annotated body-part labels. Our results demonstrate that the LLM achieves a micro-F1 score of $0.72$, indicating strong alignment with the ground truth. These results support our claim that LLMs can effectively reason about style–part alignment in a structured motion control setting.
Exploring the scope of generalization to new styles.
Although our approach leverages the reasoning capabilities of LLMs to better handle new styles, its generalization remains constrained by the representational capacity of the body-part VQ-VAE itself.
6. Conclusion
We have introduced a novel paradigm for stylized motion generation by framing it as a reasoning-composition-generation process in a body-part centric text space. By leveraging the semantic richness and reasoning ability of large language models (LLMs), our approach-SMooGPT, a fine-tuned LLM, achieves interpretable, controllable, and generalizable motion synthesis that surpasses prior methods in handling complex styles and diverse content. Extensive experiments and user studies validate the effectiveness of our method, particularly under challenging text-driven settings. We believe our work opens a new direction for motion generation by unifying natural language understanding with structured motion representation.
While our method demonstrates strong performance, the current body-part text space exhibits limited capacity in representing intricate motion dynamics. In future work, we aim to enrich this space by augmenting each body-part text with a series of keypose-based descriptions that encode spatial configurations over time. Such temporally-aware extensions will allow the model to capture motion dynamics more effectively and provide a richer representation of both content and style.
Supplemental Material
In this supplementary material, we provide additional discussions, details on dataset construction, the implementation of our method and all baselines, as well as examples of body-part text corresponding to the generated motions. Furthermore, we present extended results on pretraining and instruction tuning conducted on HumanML3D.
Appendix A More Discussions
@@
Appendix B Dataset Constructions
@@
In- and out-of-domain subset construction.
To enable a more meaningful comparison between SMooGPT and SMooDi, we split the 100STYLE dataset into in-domain and out-of-domain subsets. Since some style labels in 100STYLE inherently contain content-specific cues, we primarily follow the grouping defined in SMooDi. Additionally, we designate the following style labels as out-of-domain for evaluation: Zombie, SwingShoulders, SwingArmsRound, Swimming, Swat, Superman, Stiff, Star, ShieldedRight, and ShieldedLeft. During stylization-enhanced tuning, we include all styles except those designated as out-of-domain. For out-of-domain evaluation, we randomly sample text or motion sequences from the out-of-domain subsets as style inputs and use the model trained on the in-domain subsets to assess generalization.
Appendix C Implementation Details
@@
We leverage T5 (Chung et al., 2024) as the underlying architecture for our language model, with 12 layers in both the Transformer encoder and decoder. The model is trained with a learning rate of $2\times 10^{-4}$ during the pre-training stage and $1\times 10^{-4}$ during the instruction tuning stage, using a mini-batch size of 48 for both. Each stage consists of 300K training iterations. For the stylization-enhanced tuning, the model is trained for an additional 50K iterations. All models are trained on 2 NVIDIA H200 GPUs.
@@
To evaluate the inference efficiency of our approach, we report the average inference time of generating a single motion clip, measured in seconds, on a single NVIDIA RTX4090 GPU. Specifically, our approach takes approximately 7.99 seconds to generate a motion clip.
Appendix D Design Details of Competitors

### v2

@@ 5.1. Comparison
This task generates stylized motion from a content text and a reference style motion. To the best of our knowledge, SMooDi (Zhong et al., 2024) is the only method designed for this setting, and thus serves as our primary baseline. To evaluate generalization to unseen styles, we consider two evaluation protocols. The zero-shot (cross-dataset) setting trains both SMooGPT and SMooDi on HumanML3D and evaluates them using content texts from HumanML3D paired with style motions from 100STYLE. The held-out (intra-dataset) setting reserves $10$ styles from 100STYLE for testing, ensuring that target styles are unseen during training. To ensure a fair comparison under unseen settings, we adapt the configurations of both models. Specifically, SMooGPT directly incorporates the reference style motion into the diffusion process to avoid information loss from text translation (see supplementary), while SMooDi disables its classifier-based style guidance to avoid reliance on style supervision requiring seen styles.
The statistical and visual comparisons are presented in Tab. 2 and Fig. 6 (bottom), respectively. In the zero-shot (cross-dataset) setting, SMooGPT achieves better performance than SMooDi in all metrics other than R-Precision. Specifically, our method achieves a remarkable 78.5% improvement in style reflection. This validates that our strategy of performing composition in the body-part text space confers superior generalization capabilities for stylization. In the held-out (intra-dataset) setting, our approach maintains this advantage, surpassing SMooDi in both style expression and motion realism. Despite slightly lower R-Precision in two settings, SMooGPT consistently outperforms SMooDi on other content preservation metrics, including MM Dist and CLIP-score.
@@
Method SRA-L $\uparrow$ MM Dist $\downarrow$ R-P.(Top-3) $\uparrow$ CLIP-score $\uparrow$ FS-Ratio $\downarrow$
w/o BP 13.674 1.231 0.452 0.6324 0.0963
w/o PT 15.234 1.380 0.501 0.667 0.0958
w/o DF 16.328 1.134 0.512 0.669 0.0953
Ours 17.344 1.132 0.520 0.670 0.0832
@@ 5.2. Ablation Study
To validate the effectiveness of each module in our framework, we conduct several ablation studies on the text-guided stylization task and show the visual ablation in Fig. 7.
The importance of body-part space.
To assess the impact of incorporating the body-part text space, we conduct an ablation study using the original MotionGPT (Jiang et al., 2023) model—referred to as w/o BP—which is trained solely on global textual inputs within the same dataset. As illustrated in Tab 3, incorporating body-part information yields a 26.8% improvement in style reflection, demonstrating the advantage of fine-grained control over individual body parts. Moreover, this structured representation simultaneously enhances semantic alignment—boosting R-Precision by 15.0% and CLIP-score by 5.9%—while significantly improving motion realism by reducing the foot sliding ratio by 13.6%.
The importance of pre-training.
To assess the role of pre-training, we remove the pre-training phase (denoted as w/o PT). As shown in Tab. 3, this leads to consistent performance degradation across all metrics, including a 12.2% drop in style reflection and a substantial 21.9% increase in MM Dist. These results highlight the importance of pre-training for aligning fine-grained body-part texts with motion representations, which in turn facilitates effective post-training and improves overall performance.
The importance of the diffusion head.
To evaluate the contribution of the diffusion head, we perform an ablation study by excluding this component (denoted as w/o DF). As shown in Tab. 3, incorporating the diffusion head yields substantial performance gains in both realism and stylization. Specifically, the inclusion of the diffusion head reduces the FS-Ratio by approximately 12.7%. These improvements validate that the diffusion head plays a critical role in coordinating body-part motions, ultimately leading to more harmonious motion.
@@ 5.3. User Evaluation
Figure 8. User Evaluation Statistics. The percentage of times our approach is preferred over (a) ChatGPT+MLD, (b) ChatGPT+MotionGPT, and (c) ChatGPT+ParCo is reported. When choosing from a pair of generated motions, users are asked to evaluate three aspects - Content Preservation, Style Reflection, and Motion Realism. The higher the percentage (50% is a tie), the better our results.
We conduct a user perceptual study using pairwise comparisons on the text-guided stylization task. In each trial, participants are shown a motion content text, a motion style text, and two motion clips generated by SMooGPT and a competing method (ChatGPT+MLD, ChatGPT+MotionGPT, or ChatGPT+ParCo), and are asked to evaluate content preservation, style reflection, and motion realism (Fig. 8). We invited $75$ participants in total, and Fig. 8 reports the percentage of times SMooGPT is preferred over competing methods.
Overall, SMooGPT is consistently favored across all three aspects. Specifically, it achieves preference rates of $81.3\%$, $81.7\%$, and $80.7\%$ against ChatGPT+MotionGPT, remains above $72\%$ when compared with ChatGPT+MLD, and obtains similarly strong results against ChatGPT+ParCo ($79.57\%$, $75.1\%$, and $79.53\%$). Overall, the user perceptual study validates our superior performance and is consistent with observations in the comparison.
[…]
@@ 5.4. Discussions
Fueled by the reasoning capabilities of LLMs, our approach acts as an effective composer that reconciles motion content and style texts in a conflict-free manner, which is critical when the two impose contradictory constraints on specific body parts. Fig. 9 shows a representative example: the motion content “throw” requires one arm to stay close to the body, while the “Aeroplane” style demands both arms be lifted and extended outward. Our method resolves this conflict by selectively borrowing one arm configuration from each source, producing a composed motion where one arm performs the throwing action and the other adopts the aeroplane pose. In contrast, SMooDi, which takes style motion rather than style text as input, fails to coherently satisfy both content and style.
@@
6. Conclusion
We propose a new paradigm for stylized motion generation by framing it as a reasoning–composition–generation process in a body-part-centric text space. By leveraging the semantic richness and reasoning ability of large language models (LLMs), our SMooGPT, a fine-tuned LLM, achieves interpretable, controllable, and generalizable stylized motion synthesis that surpasses prior methods in handling novel style. Extensive experiments and user studies validate the effectiveness of our method. We believe our work opens a new direction for motion generation by unifying natural language understanding with structured motion representation.
Supplemental Material
In this supplementary material, we provide additional details that support and complement the main paper. Specifically, we include detailed descriptions of dataset construction, implementation details of our method and all baselines, and extended examples of body-part text representations corresponding to the generated motions. We provide a supplemental video, which we encourage the reviewers to watch since motion is critical in our results, and this is hard to convey in a static document.
Appendix A More Discussions
Semantic composition vs. latent disentanglement.
Unlike prior methods that rely on explicit feature disentanglement constraints to separate style from content—a process often fraught with optimization instability—we propose a paradigm shift towards implicit semantic composition. By projecting both motion content and style into a unified, interpretable body-part text space, we bypass the need for complex losses or specialized architectures to enforce disentanglement. Our framework leverages the reasoning capabilities of LLMs to merge these attributes at the linguistic level, ensuring that style and content are naturally integrated into conflict-free body-part instructions before generation.s
LLM-driven motion priors and body-part decomposition
The core assumption of SMooGPT is that LLMs possess inherent knowledge of human motion and motion styles, alongside the capability to decompose complex movements into body-part-level components. Based on this, we establish a reason-compose-generation pipeline that leverages LLMs for semantic reasoning. Recent research (Li et al., 2025c) provides strong empirical evidence for this approach. Their findings confirm that LLMs excel at high-level movement planning and can accurately deconstruct complex instructions into part-specific descriptions. This directly validates our ”part-based decomposition” strategy. Furthermore, they demonstrate that LLMs exhibit robust ”motion commonsense” for creative and culturally specific styles that are often underrepresented in traditional mocap datasets. This finding provides critical support for SMooGPT’s assumption regarding LLMs’ inherent knowledge of motion styles.
Style and Body-Part Alignment Capability.
To assess whether large language models can reason about the relationship between motion styles and the body parts they primarily affect, we conduct a quantitative evaluation on the 100STYLE dataset. Since the original dataset does not provide explicit body-part annotations for styles, we manually annotate each style by inspecting representative motion sequences and identifying the body parts that exhibit the most prominent stylistic variations. Each style is annotated with one or more body-part labels, forming a multi-label ground-truth set.
For evaluation, we prompt GPT using only the global style name and its textual description, without providing any motion data, and ask it to predict the relevant body parts from a predefined list. We then compare the predicted body-part sets with the human-annotated ground truth using the micro-F1 score under a multi-label classification setting. The resulting micro-F1 score of $0.72$ indicates a strong alignment between the LLM predictions and human annotations, supporting our claim that LLMs possess a meaningful prior for reasoning about style–body-part relationships in structured motion control.
The coordination of body-part motion.
SMooGPT utilizes two coordination mechanisms: (1) implicit coordination, occurring during the LLM-driven translation from body-part text to motion tokens, and (2) explicit coordination, in which the diffusion head directly models the dependencies between body-part motion tokens.
Appendix B Instruction Templates for Post-training
This section provides the instruction templates used for post-training with instruction following, as described in Sec. 4.3 of the main paper.
Global-to-Body-Part Decomposition.
This instruction trains the model to decompose a global motion description into fine-grained body-part-centric texts, where each text describes the localized motion behavior of a specific body part.
Input: Global motion text $\mathbf{T}_{g}$>
Content–Style Composition.
This instruction trains the model to compose motion content and motion style descriptions in the body-part space, resolving potential conflicts between them to produce a unified and coherent description.
Appendix C Dataset Constructions
@@
Held-out style subset construction.
To enable a fair comparison between SMooGPT and SMooDi under the held-out setting, we partition the 100STYLE dataset into seen and held-out style subsets. Since some style labels in 100STYLE inherently encode content-specific cues, we primarily follow the style grouping protocol adopted in SMooDi to avoid content leakage. Specifically, we designate the following styles as held-out and exclude them entirely from training: Zombie, SwingShoulders, SwingArmsRound, Swimming, Swat, Superman, Stiff, Star, ShieldedRight, and ShieldedLeft. During stylization-enhanced tuning, the model is trained only on the remaining seen styles and is evaluated on the held-out styles to assess generalization.
For out-of-domain evaluation, we randomly sample text or motion sequences from the out-of-domain subsets as style inputs and use the model trained on the in-domain subsets to assess generalization.
Appendix D Implementation Details
Evaluation Metrics.
Given that the standard evaluator (Guo et al., 2022a) can bias embeddings toward text alignment at the expense of motion fidelity, we instead utilize TMR (Petrovich et al., 2023) to compute MM-Dist and R-Precision. This choice is consistent with recent protocols (Hwang et al., 2025) to ensure a more robust assessment of content preservation.
@@ Appendix D Implementation Details
Diffusion-based Motion Refinement
While LLMs can generate conflict-free body-part texts and translate them into motion tokens, this process operates sequentially in a discrete space and lacks explicit modeling of continuous coordination across body parts. To bridge this gap, we introduce a lightweight diffusion head preceding the decoder. Let $\mathcal{C}_{in}=\{c_{p}\}_{p=1}^{P}$ denote the set of independent body-part motion latents predicted by the LLM. These inputs are first projected and composed into a unified continuous representation $\bm{z}_{\text{cond}}=\Psi(\mathcal{C}_{\text{in}})$ using a composition function $\Psi$ implemented as a set of MLPs followed by a concatenation operation.
The core of this module is the denoiser $\epsilon_{\theta}$ parameterized by $\theta$. Architecturally, we implement $\epsilon_{\theta}$ as a set of MLPs followed by a concatenation operation. The diffusion process operates on a sequence of unified motion latents $\{\bm{z}_{t}\}_{t=0}^{T}$, starting from Gaussian noise $\bm{z}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$. At each denoising step $t$, the model predicts the noise component to refine the latent:
(4) $\epsilon_{t}=\epsilon_{\theta}(\bm{z}_{t},t,\bm{z}_{\text{cond}}).$
We optimize the model using the standard noise reconstruction loss:
(5) $\mathcal{L}_{\text{recon}}=\mathbb{E}_{\epsilon,\bm{z}_{0},t}\left[\left\|\epsilon_{\theta}(\bm{z}_{t},t,\bm{z}_{\text{cond}})-\epsilon\right\|^{2}_{2}\right],$
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ represents the ground-truth noise added to the clean latent $\bm{z}_{0}$.
After $T$ iterative denoising steps, a clean and coordinated unified latent $\bm{z}_{0}$ is obtained. Crucially, applying the inverse operation $\Psi^{-1}$ decomposes $\bm{z}_{0}$ back into body-part representations, resulting in $\{\hat{c}_{p}\}_{p=1}^{P}$. Finally, these refined body-part latents are decoded by $\mathcal{D}$ into the final motion sequence $\bm{m}\in\mathbb{R}^{N\times H}$.
Motion-guided stylization.
In this task, we aim to synthesize stylized motion conditioned on both content text and a style motion sequence. A straightforward approach involves using SMooGPT to first translate the style sequence into body-part textual descriptions, compose them with the content text, and subsequently generate the output. However, projecting high-frequency style motion into the discrete text domain inevitably results in a loss of fine-grained dynamic details. To mitigate this, drawing inspiration from SMooDi (Zhong et al., 2024), we employ a lightweight adapter to inject style guidance into the diffusion head. Specifically, the denoising step is modified to condition on the style embedding. Let $\bm{s}$ denote the style features extracted from the reference motion $\bm{m}_{r}$. The noise prediction network $\epsilon_{\theta}$ is updated to:
(6) ${\epsilon}_{t}=\epsilon_{\theta}(\bm{z}_{t},t,\bm{z}_{\text{cond}},\bm{s}).$
The standard noise reconstruction loss is updated to:
(7) $\mathcal{L}_{\text{recon}}=\mathbb{E}_{\epsilon,\bm{z}_{0},t}\left[\left\|\epsilon_{\theta^{\prime}}(\bm{z}_{t},t,\bm{z}_{\text{cond}},\bm{s})-\epsilon\right\|^{2}_{2}\right].$
In the zero-shot setting, models are trained solely on HumanML3D. Conversely, in the intra-dataset setting, training is performed on HumanML3D augmented with the seen subset of 100STYLE.
@@
We leverage T5 (Chung et al., 2024) as the underlying architecture for our language model, with 12 layers in both the Transformer encoder and decoder. The model is trained with a learning rate of $2\times 10^{-4}$ during the pre-training stage and $1\times 10^{-4}$ during the instruction tuning stage, using a mini-batch size of 48 for both. Each stage consists of 300K training iterations. To prevent the model from over-relying on the global text or specific body parts, we randomly mask out either the global text or one to two body-part texts during training. For the stylization-enhanced tuning, the model is trained for an additional 50K iterations. All models are trained on 2 NVIDIA H200 GPUs.
@@
To evaluate the inference efficiency of our approach, we report the average inference time of generating a single motion clip, measured in seconds, on a single NVIDIA RTX4090 GPU. Specifically, our approach takes approximately 25.761 seconds to generate a motion clip.
Appendix E Design Details of Competitors